Solve the following equation:
(x + 4) (x + 3) = 6

Found 2 solutions by nerdybill, padmameesala:
Answer by nerdybill(7384)

(x + 4) (x + 3) = 6
First, expand the left with FOIL:
x^2 + 3x + 4x + 12 = 6
x^2 + 7x + 12 = 6
x^2 + 7x + 6 = 0
Now, factoring the left side:
(x+1)(x+6) = 0
.
x = {-1, -6}

Answer by padmameesala(42)

(x + 4)(x + 3) = 6
x^2 + 3x + 4x + 12 = 6
x^2 + 7x + 12 - 6 = 6 - 6
x^2 + 7x + 6 = 0
x^2 + 6x + x + 6 = 0
x(x + 6) + 1(x + 6) = 0
(x + 1)(x + 6) = 0
x = -1,-6